How do you change these equations to slope y-intercept form? -3x-y=-1 and 2x+4y=-16 thanks
Answer by rfer(16322)

y=mx+b
-------------
-3x-y=-1
-y=3x-1
y=-3x+1
--------------
2x+4y=-16
4y=-2x-16
y=-1/2x-4